# Challenges and Opportunities for Visual Analytics in Jurisprudence

### v2

@@ 7.4 Future Work and Generalization
To address the limited evaluation, we plan to implement the proposed VA workflow in preparation for a systematic evaluation. In this context, corresponding user studies should involve controlled experiments conducted with legal scholars. These experiments should compare the efficacy and usability of traditional UI, which utilizes syntactic and keyword-based search interfaces, with the proposed VA-enhanced workflow. Additionally, the experiments should assess the reduction in cognitive load experienced when utilizing interactive visualizations in legal research tasks.
@@
Our work focuses primarily on the German civil law system. While the civil law system works similarly, countries sometimes have substantial differences regarding style, structure, and methodology. For example, EU legislation and German law differ in articulating their goals. In EU legislation, goals are clearly outlined, while German law lacks such explicit statements, sometimes to be found in the law justifications. Additionally, the two systems vary in their teleological and systematic interpretations. There are also differences in the integration of certain aspects of case law into EU legislation as a supranational legal system. Future work should explore the particularities of extending our approach to other legislative systems and, in particular, to case law, with more considerable differences. Regardless, we expect that our distilled three-phase workflow (refer to subsection 5.1) still applies. The three key areas of VA also benefit jurisprudence in these legal systems due to their generalizability.
Unfortunately, open legal datasets necessary for training and fine-tuning are limited and need more diversity. While several German legal datasets include fundamental law and some court rulings (refer to Table 1), law commentary, interpretation, and the vast majority of legal cases and statements of claims are rarely available. However, access to law commentary is fundamental for the legal reasoning process [4, p. 254]. In the future, more work must be done to offer more structured open legal data. One possible direction forward is through open data initiatives, digitalizing legal systems, or scanning older literature that is automatically transformed into legal archives with semantic LLM-based methods alongside reforms in copyright legislation and fair use.
While we have presented the textual analysis with a strong focus on jurisprudence, some methods and solutions exhibit substantial overlap with workflows in related, text-based research domains [62, 60]. These workflows share core elements of data navigation, knowledge representation, and analytical reasoning, as outlined in section 6. Consequently, our VA workflow (refer to subsection 7.1) can be transferred onto these usage scenarios, albeit with domain-specific adaptations. For example, the types of analytical artifacts must align with the target discipline. In the context of scientific literature, scholars writing a survey about visualization types might use the workflow to stash evidence regarding papers covering specific visualization techniques. In this case, a primitive artifact could represent an individual visualization figure, while a compound artifact could encapsulate a collection of papers discussing the same visualization type. This adaptability stems from the workflow’s abstraction into the broader sense-making loop [93], as described by Pirolli and Card. Implementations of this loop are evident in various domains, including healthcare [45] and sports [122]. For instance, the VA system proposed by Chung et al. for analyzing rugby performance [122] uses a projection of game events to facilitate data navigation. This approach parallels the projection of medical texts employed by KMTLabeler [111]. Additionally, both systems enable domain experts to incorporate their tacit knowledge through weight modifications during analysis [122, 111].
Thus, our VA workflow can provide a guiding framework for visualization researchers across diverse domains, where tacit knowledge remains a driving force for analytical reasoning and decision-making. Several existing systems externalize tacit domain knowledge as explicit rules, often as tables [122]. While functional, these representations may appear overly generic and disconnected from the specific workflows of individual domains. Therefore, future research should prioritize the development of domain-specific metaphors for externalizing tacit knowledge [48].
@@ 8 Conclusion
In this work, we explore the integration of Visual Analytics (VA) into jurisprudence, addressing the unique challenges of legal texts’ complex, formal, and hierarchical nature. Through semi-structured interviews with legal experts, we identify a typical workflow in jurisprudence and analyze its commonalities with workflows in other disciplines. This analysis reveals an abstraction in the sense-making loop, emphasizing the need for VA systems that integrate data navigation, knowledge representation, and analytical reasoning within a unified design. To address these gaps, we propose a human-is-the-loop VA design that leverages legal scholars’ tacit knowledge, enhancing their ability to navigate, interpret, and reason with legal documents. We critically discuss the necessary transformations from a traditional UI commonly employed to a VA-based interface, alongside its challenges. Our proposed approach facilitates informed legal reasoning by combining human expertise with machine intelligence. It provides a scalable framework for other text-intensive domains and the blueprint for future development of such systems. Our findings extend beyond jurisprudence, offering valuable insights into the design of VA systems for related fields. This work lays a foundation for future advancements in knowledge-assisted VA systems by addressing the interplay between tacit knowledge and sense-making processes.
Acknowledgements / Funding
This work has been funded by the Federal Ministry of Education and Research (BMBF) in VIKING (13N16242), the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) in RATIO-CUEPAQ (455910360), and under Germany’s Excellence Strategy – EXC 2117 – 422037984.
@@
Statements and Declarations
Conflict of interest — There is no conflict of interest or competing interests from any of the authors.
Consent to participate — All participants were informed and gave their consent to participate in the study. Participants were informed about the use of the questionnaire/interview data for an article.
Consent for publication — All participants gave their consent to publish the data collected in the surveys and related to their participation in the study. We do not publish any identifying information about the participants.

### v3

@@ 7.4 Future Work and Generalization
To address the limited evaluation, we plan to implement the proposed VA workflow for a systematic evaluation. In this context, corresponding user studies should involve controlled experiments comparing the efficacy and usability. Additionally, the experiments should assess the cognitive load experienced when utilizing interactive visualizations.
@@
Our work is scoped to the German civil law system, with qualitative insights derived from nine legal experts representing various legal disciplines within this system. Consequently, the proposed VA design fits this context but offers a methodological framework that could be replicated in other legal environments within the civil law system. While these environments work similarly, countries sometimes have substantial differences regarding style, structure, and methodology. For example, EU legislation and German law differ in articulating their goals. In EU legislation, goals are clearly outlined, while German law lacks such explicit statements, sometimes to be found in the law justifications. Additionally, the two systems vary in their teleological and systematic interpretations. There are also differences in the integration of certain aspects of case law into EU legislation as a supranational legal system. Future work should explore the particularities of extending our approach to other legislative systems and, in particular, to case law, with considerable differences. For replication, researchers would need to interview domain experts familiar with the legal workflows in that system to validate the adaptation of our three-phase workflow. Additionally, the researchers would need to have visualization expertise to verify that the adaption still aligns with the proposed VA design.
While we have presented the textual analysis with a strong focus on jurisprudence, some methods and solutions exhibit substantial overlap with workflows in related, text-based research domains [zengVIStoryInteractiveStoryboard2021, tianLitVisVisualAnalytics2023]. These workflows share core elements of data navigation, knowledge representation, and analytical reasoning, as outlined in 6. Consequently, our VA workflow (see 7.1) can be transferred onto these usage scenarios, albeit with domain-specific adaptations. We suppose that the workflow phases and the visualization techniques used to support data navigation, knowledge representation, and analytical reasoning are domain-agnostic, while the underlying data, its hierarchical structure, and details of the reasoning processes are domain-specific. Hence, similar to jurisprudence, the domain transferred to must offer accessible data sets that impose a hierarchical structure on the domain semantics.
Overall, our VA workflow can provide insights and a starting ground for visualization researchers in domains where tacit knowledge remains a driving force for analytical reasoning and decision-making.
@@ 8 Conclusion
In this work, we explore the integration of Visual Analytics (VA) into jurisprudence, addressing the unique challenges of legal texts’ complex, formal, and hierarchical nature. Through semi-structured interviews with legal experts, we identify a typical workflow in jurisprudence and analyze its commonalities with workflows in other disciplines. This analysis reveals an abstraction in the sense-making loop, emphasizing the need for VA systems that integrate data navigation, knowledge representation, and analytical reasoning within a unified design. To address these gaps, we propose a human-is-the-loop VA design that leverages legal scholars’ tacit knowledge, enhancing their ability to navigate, interpret, and reason with legal documents. We critically discuss the necessary transformations from a traditional user interface commonly employed to a VA-based interface, alongside its challenges. Our proposed approach facilitates informed legal reasoning by combining human expertise with machine intelligence. It provides a scalable framework for other text-intensive domains and the blueprint for future development of such systems. Our findings extend beyond jurisprudence, offering valuable insights into the design of VA systems for related fields. This work lays a foundation for future advancements in knowledge-assisted VA systems by addressing the interplay between tacit knowledge and sense-making processes.
@@
Funding
This work has been funded by the Federal Ministry of Education and Research (BMBF) in VIKING (13N16242), the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) in RATIO-CUEPAQ (455910360), and under Germany’s Excellence Strategy - EXC 2117 - 422037984.
Declarations
Conflict of interest
There is no conflict of interest or competing interests from any of the authors.
Consent to participate
All participants were informed and gave their consent to participate in the study. Participants were informed about the use of the questionnaire/interview data for an article.
Consent for publication
All participants gave their consent to publish the data collected in the surveys and related to their participation in the study. We do not publish any identifying information about the participants.